A coupon bond pays annual interest, has a par value of $1,000, matures in 5 (five) years, has a coupon
rate of 7.45%, and has a yield to maturity of 8.82%. The current yield on this bond is ________%. (2 decimal place)
Answer #1

Annual coupon=1000*7.45%=$74.5

Current price=Annual coupon*Present value of annuity factor(8.82%,5)+1000*Present value of discounting factor(8.82%,5)

=74.5*3.9078857+1000*0.655324481

=$946.461966(Approx)

Current yield=Annual coupon/Current value

=74.5/946.461966

=7.87%(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=74.5[1-(1.0882)^-5]/0.0882

=74.5*3.9078857

2.Present value of discounting factor=1000/1.0882^5

=1000*0.655324481